John was given a residence in the current year. At the time of the gift, the residence had a fair market value of $250,000, and its adjusted basis to the donor was $130,000. The donor paid a tax of $12,000 on the taxable gift of $236,000. What is John’s basis for gain?

a- $130,000.

b- $136,102.

c- $142,000.

d- $250,000.

e- $236,000.
Answer #1

Answer is option B

b- $136,102.

John’s basis for gain = adjusted basis + proportional gift tax = 130000 + ((250000-130000)/236000*12000) = 130000+6102= $136102